OpenStudy (anonymous):

What is the point-slope form of the equation of a line that passes through the point (1, 4) and has a slope of −3?

OpenStudy (anonymous):

Do you know how to start the problem?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

well point slope intercept form is y=mx+b You have x and y. is -3 m or b?

OpenStudy (anonymous):

I'm not typing don't believe that. -3 is m. Your equation is 4=-3(1)+b you solve the equation to find b

OpenStudy (triciaal):

point -slope y - y1 = m(x - x1) slope-intercept y = m x + b

OpenStudy (triciaal):

the question says point-slope m = slope = -3 and point given (x1, y1) = (1, 4) any point (x, y) y - 4 = -3 ( x - 1)

OpenStudy (triciaal):

@AJLYAMBER96 do you see now there is no "point slope intercept form" ?

OpenStudy (anonymous):

that was my bad. Yes I see what you mean. I was relating the two equations.